In attendance: J. Nelson (notetaker), R. Chestnut, R. Hall, J. Rock
New agenda items:
Janice will work with Bob
Oracle is the bottleneck. Entries are indexed and partitioned. Bob will work with Ian MacGregor.
SLED FDBK through SCP slow feedback
Ed Miller comes back 2/9. Joe will be back on Thursday.
SIS waveforms archived every hour (plus on change)
We do want this and around early April
SISDTZ signal updates without archiving
Janice will double check again.
Text string PV for shift taker name
Judy hooked us up. And it’s beautiful.
Old agenda items:
List of phase 2a signals
Judy and Ron are working on it. Still need channel assignment information from Steve Smith (ADC channel info, etc). Right now, the only ADC channels being used are 0 & 1 (twt5:inp:fwd and twt6:inp:fwd). If these need to be lobotomized, please let Janice know and she/Joe can move wires around.
macnair module support of aes
Ron is working on this. Progress from Dusatko. Will check ‘em into tr09 and try again. (Joe &/or Linda may be called.)
Summary PVs for plc panels? About to finish…. Will test after the meeting. Yeah. Then will add to displays
Nice email from Alan Hill. Janice has instructions for what to try when installed and ready to go.
Gtw04 getting bigger brain and faster hamster tomorrow during DBInstall. All relevant nlcta & 8pack processes will be run on gtw04. Gtw5 will be used only for fall back. (You can still log into gtw05 and access PVs.)
This is ongoing. Ops should expect just a few small interruptions in the future.
TR06 is coming…